# VoxGuard: Evaluating user and attribute privacy in speech via Membership Inference Attacks

### v1

@@ 1 Introduction
Speech recognition systems raise significant privacy concerns because spoken data reveals both linguistic content and biometric identity. Regulatory frameworks, such as GDPR [1], explicitly classify voice as personal data and, in some cases, as biometric data. This dual nature of voice raises an important question: what types of sensitive information can be inferred from speech? On the one hand, the linguistic channel conveys the literal content of what is said. On the other hand, the paralinguistic channel, such as tone, pitch, and timbre, forms a distinctive "voiceprint" that can uniquely identify a speaker [2]. Apart from speaker identification, voice also encodes a wide range of private attributes such as gender, accents, emotional state, and even health conditions [3, 4, 5], which individuals may prefer to keep confidential. In this work, we direct part of our analysis toward these personal attributes.
@@
This paper uses viewpoints from Differential Privacy, particularly on membership inference, to evaluate speech privacy. We introduce VoxGuard, an evaluation framework consisting of two complementary notions: User Privacy corresponding to the traditional notion of preventing speaker re-identification and Attribute Privacy quantifying the risk of inferring speaker traits (e.g., gender, accent). For each, we define worst- and average-case notions aligned with DP-style indistinguishability and evaluate at low-FPR operating regions to expose worst-case leakage. Empirically, we find that EER can diverge from measuring privacy at low-FPR regimes and we provide upper-bounds for the attacks.
@@ 3.3 Membership Inference and Low-FPR Metrics
Figure 1: Illustration of the four VoxGuard evaluation tasks. User-Privacy-Strict: adversary compares same-text pairs to decide if two samples belong to the same speaker (identity as the varying factor). User-Privacy-Relaxed: adversary compares arbitrary texts from two speakers, reflecting the general case where content varies. Attr-Privacy-Strict: pairs are matched on all attributes except one, isolating the effect of a single attribute. Attr-Privacy-Relaxed: pairs differ in one attribute without controlling for others, modeling the general case where multiple traits vary.
@@ 5.2 Attribute Privacy
Figure 2: Comparing ROC curves in log–log scale reveals severe gaps between User Privacy attacks. While average-case EERs range only from 20–47%, their performance at low FPR differs drastically. FT–Max at $L=200$ achieves TPR >>50% at $\text{FPR}=10^{-4}$, whereas PT–Avg remains near random. This shows that reporting only EER obscures strong privacy leaks. Max scoring and fine-tuned attackers yield orders-of-magnitude stronger attacks in the low-FPR regime.

### v2

@@ Abstract
Index Terms—  Privacy, ASV, Speech, Differential Privacy, Membership Inference
@@ 1 Introduction
Speech recognition systems raise significant privacy concerns because spoken data reveals both linguistic content and biometric identity. Regulatory frameworks, such as GDPR [1], explicitly classify voice as personal data and, in some cases, as biometric data. This dual nature of voice raises an important question: what types of sensitive information can be inferred from speech? On the one hand, the linguistic channel conveys the literal content of what is said. On the other hand, the paralinguistic channel, such as tone, pitch, and timbre, forms a distinctive ”voiceprint” that can uniquely identify a speaker [2]. Apart from speaker identification, voice also encodes a wide range of private attributes such as gender, accents, emotional state, and even health conditions [3, 4, 5], which individuals may prefer to keep confidential. In this work, we direct part of our analysis toward these personal attributes.
@@
This paper uses viewpoints from Differential Privacy, particularly on membership inference, to evaluate speech privacy. We introduce VoxGuard, an evaluation framework consisting of two complementary notions: User Privacy corresponding to the traditional notion of preventing speaker re-identification and Attribute Privacy quantifying the risk of inferring speaker traits (e.g., gender, accent). For each, we define strict- and relaxed-case notions aligned with DP-style indistinguishability and evaluate at low-FPR operating regions to expose worst-case leakage. Empirically, we find that EER can diverge from measuring privacy at low-FPR regimes and we provide upper-bounds for the attacks.
@@ 3.3 Membership Inference and Low-FPR Metrics
Fig. 1: Illustration of the four VoxGuard evaluation tasks. User-Privacy-Strict: adversary compares same-text pairs to decide if two samples belong to the same speaker (identity as the varying factor). User-Privacy-Relaxed: adversary compares arbitrary texts from two speakers, reflecting the general case where content varies. Attr-Privacy-Strict: pairs are matched on all attributes except one, isolating the effect of a single attribute. Attr-Privacy-Relaxed: pairs differ in one attribute without controlling for others, modeling the general case where multiple traits vary.
@@ 5.2 Attribute Privacy
Fig. 2: Comparing ROC curves in log–log scale reveals severe gaps between User Privacy attacks. While average-case EERs range only from 20–47%, their performance at low FPR differs drastically. FT–Max at $L=200$ achieves TPR >>50% at $\text{FPR}=10^{-4}$, whereas PT–Avg remains near random. This shows that reporting only EER obscures strong privacy leaks. Max scoring and fine-tuned attackers yield orders-of-magnitude stronger attacks in the low-FPR regime.
@@ 6 Conclusion
7 Acknowledgements
This work was supported by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the ARTS Program under contract D20232308110001 and the HIATUS Program under contract #202222072200006. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.